Atsie (atsie):

Can someone give me a few idea's for a real life situation, and help me through this question please? I'd really really appreciate it! ^.^

Atsie (atsie):



Atsie (atsie):

Here, if the picture isn't revealing it fully, I can copy it here as well.... "Describe a real-life situation that can be represented using a relation and discuss how one of the quantities in the relation depends on the other. Then give an example of such a relation in three different ways."

Atsie (atsie):

@mathmale @welshfella :)

OpenStudy (tanner23456):

I'm not completely sure on what it's asking here. I would assume it's referring to a standard y = mx+b type problem. Where, for every x input, you get a different why output. For example, a person makes a base payment of let's say $100 and then they have to pay $5 every month. Y = 5x + 100 If you put in 3 months for x, you'd get an output of $115. Is this even close to what you're looking for? You could also look at the directly related and inversely related equations. They follow the same principle. y = kx y = k/x

Atsie (atsie):

Woah, I was definitely not taught that. Its funny, the information I'm taught is usually very lacking at my school. The whole point of the chapter that I'm given is to represent relations and determine domain, range, and inverse relations which I learned apparently is about switching the coordinates in each order pair. This whole question seems related to those type of relations, so I suppose the principles of y = kx or y = k /x does me no good at the moment.

Atsie (atsie):

But, with that being said, what if we used the "real-life" situation that you just came up with. Talking about a person making a base payment of a $100, and paying @5 each month. Wouldn't we have to come up with different number's in order to compare and contrast relations like its asking? "discuss how one of the quantities in the relation depends on the other" We probably would need to add more to the situation, right? If so, what?

Atsie (atsie):

Oops. $5 I meant. :)

OpenStudy (tanner23456):

Yeah, sorry, this question is just really general. The idea of the equation y = 5x + 100 is that the value of y depends on the x value. So, whatever number you pick for x, it changes the y value. Therefore the y value depends on the x value and vice versa

Atsie (atsie):

So what if I picked 10 for the x value. 10 being...I don't something random like tax. (Overall that doesn't make sense, because more than likely a person paying base payment in that type of situation wouldn't work, but just generally speaking for now)